Find 3 consecutive even numbers where the product of the smaller two numbers is 52 less than the square of the largest number.
Answer by mananth(12270)

let the consequtive even numbers be x,x+2 , x+4
Product of smaller two = x(x+2)
.
square of the largest = (x+4)^2
.
x(x+2)=(x+4)^2 -52
x^2+2x= x^2+8x+16-52
6x=36
x= 6
The numbers are 6,8,10